OpenStudy (elitekage):

Write an equation of a line that passes through the point (3, 2) and is parallel to the line y = 3x − 4.

MARC (marc_d):

gradient,m=3

MARC (marc_d):

given that the information it must pass throught point(3,2) and must be parallel to line y=3x-4(which means that the both lines must hv the same gradient)

MARC (marc_d):

step 1: use the slope-intercept form y=mx+c where m=gradient c=y-intercept to find the value of c(y-intercept)

MARC (marc_d):

now,u hv the value of m which is 3 and a point(3,2)=(x,y) plug all the values to find the value of c(y-intercept)

MARC (marc_d):

\[\huge{y=mx+c}\]

MARC (marc_d):

\[\huge{2=3(3)+c}\]

MARC (marc_d):

\[\huge{2=9+c}\]

MARC (marc_d):

\[\huge{c=-7}\]

MARC (marc_d):

now,u hv the value of c..lastly,u can start to form an equation..

MARC (marc_d):

\[\huge{y=\color{red}{m}x+\color{red}{c}}\]

MARC (marc_d):

\[\huge{y=\color{red}{(3)}{x}+\color{red}{(-7)}}\]

OpenStudy (elitekage):

thanks

MARC (marc_d):

yw! @EliteKage